# Boundary rigidity of systolic and Helly complexes

### v1

@@ Acknowledgements
J.C. was partially supported by ANR project DUCAT (ANR-20-CE48-0006). The work presented here was partially supported by the Carlsberg Foundation, grant CF23-1225.
@@ 2.1. Graphs and simplicial complexes
All graphs $G=(V(G),E(G))$ considered in this paper are finite, undirected, connected, and contain neither multiple edges, nor loops. For two distinct vertices $v,w\in V(G)$ we write $v\sim w$ when there is an edge connecting $v$ with $w$. The subgraph of $G$ induced by a subset $A\subseteq V(G)$ is the graph $G[A]=(A,E^{\prime})$ such that $uv\in E^{\prime}$ if and only if $uv\in E(G)$. The distance $d(u,v)=d_{G}(u,v)$ between two vertices $u$ and $v$ of a graph $G$ is the length of a shortest $(u,v)$–path. An induced subgraph $H=G[A]$ of a graph $G$ is an isometric subgraph of $G$ if $d_{H}(u,v)=d_{G}(u,v)$ for any two vertices $u,v\in A$. The interval $I(u,v)$ between $u$ and $v$ consists of all vertices on shortest $(u,v)$–paths, that is, of all vertices (metrically) between $u$ and $v$: $I(u,v)=\{x\in V(G):d(u,x)+d(x,v)=d(u,v)\}$. An induced subgraph of $G$ is called convex if it includes the interval of $G$ between any pair of its vertices. For a vertex $v$ of $G$ and an integer $r\geq 1$, we will denote by $B_{r}(v)$ the ball in $G$ (and the subgraph induced by this ball) of radius $r$ centered at $v$, i.e., $B_{r}(v)=\{x\in V(G):d(v,x)\leq r\}.$ More generally, the $r$ball around a set $A\subseteq V(G)$ is the set (or the subgraph induced by) $B_{r}(A)=\{v\in V(G):d(v,A)\leq r\},$ where $d(v,A)=\mbox{min}\{d(v,x):x\in A\}$. Similarly, the sphere in $G$ centered at $v$ of radius $r$ is the set (or the subgraph induced by) $S_{r}(v)=\{x\in V(G):d(v,x)=r\}$. As usual, $N(v)=S_{1}(v)$ denotes the set of neighbors of a vertex $v$ in $G$, and $N[v]=B_{1}(v)$. The link of $v\in V(G)$ is the subgraph of $G$ induced by $N(v)$.
@@
An abstract simplicial complex $X$ on a finite set $V$ is a set of nonempty subsets of $V$, called simplices, such that any nonempty subset of a simplex is also a simplex. The dimension of a simplex $\sigma$ is $|\sigma|-1$ and the dimension of $X$ is the largest dimension of a simplex of $X$. If $\sigma$ is a simplex of dimension $k$, then all simplices of $X$ contained in $\sigma$ and having dimension $k-1$ are called the facets of $\sigma$. Simplices of $X$ are called maximal if they are maximal by inclusion, and non-maximal otherwise. The 2-dimensional simplices of $X$ are called triangles, and if a triangle is spanned by vertices $u$, $v$ and $w$, then we denote it by $uvw$.
@@
For an abstract simplicial complex $X$, the $k$–skeleton $X^{(k)}$ of $X$ consists of all simplices of dimension at most $k$. We use the notations $V(X)=X^{(0)}$ for the set of vertices (0-simplices) of $X$ and $G(X):=X^{(1)}$ for the 1-skeleton of $X$; the graph $G(X)$ will be always endowed with the standard graph distance $d_{G}$. The clique complex of a graph $G$ is the abstract simplicial complex having the cliques (i.e., complete subgraphs) of $G$ as simplices. A simplicial complex $X$ is a flag simplicial complex if $X$ is the clique complex of its $1$–skeleton. Note that flag complexes are uniquely determined by their 1-skeleton.
@@
Each abstract simplicial complex $X$ gives rise to a geometric simplicial complex $||X||$ obtained by replacing each simplex $\sigma$ of $X$ by an Euclidean simplex $||{\sigma}||$ such that gluing maps between simplices are isometries (see the book of Hatcher [22]). Vice-versa, the set of simplices of a geometric simplicial complex define an abstract simplicial complex. Although the definitions of boundary and boundary rigidity presented in the next subsection are motivated by geometric simplicial complexes, all these notions can be defined in equivalent way for abstract simplicial complexes. Therefore, in the rest of the paper by a “simplicial complex” we will mean a finite abstract simplicial complex.
@@ 2.3. Bridged and Helly graphs
    Quadrangle Condition: for $v,x,y,u\in V$ with $d(v,x)=d(v,y)=k=d(v,u)-1$ and $u\sim x,y$, $x\nsim y$, $\exists z\in B_{k-1}(v)$, $z\sim x,y$.
@@
Let $G$ be a Helly graph, and $z$ and $u$ be two vertices of $G$ with $k=d(z,u)\geq 2$. Then there exists a vertex $v$ which has distance $k-1$ to $z$ and is adjacent to $u$ and to all neighbors $w$ of $u$ with $d(z,w)\leq k$.
@@ 3. Boundary rigidity of systolic complexes
In this case, we set $X^{\prime}=X\setminus\{u\}$, and $S(X^{\prime})=S(X)\setminus\{u\}\cup\{v\}$. To derive the distance matrix $D^{\prime}$ of $S(X^{\prime})$ from the distance matrix $D$ of $S(X)$, we need to compute the distances $d(v,x)$ from $v$ to all vertices $x$ of $S(X^{\prime})$. Pick any $x\in S(X^{\prime})$. If $x\in N(u)$, then $d(v,x)=1$ since $v$ dominates $u$. Otherwise, let $k=d(x,u)$. If there exists $w\in S(X^{\prime})\cap N(u)$ with $d(x,w)=k+1$, then we can set $d(x,v)=k$. Indeed, $u$ belongs to a shortest $(w,x)$-path $P$ of $G$. If $y$ is the neighbor of $u$ in $P$, then $v\sim w,y$ because $v$ dominates $u$ in $G$. Consequently, $v$ is on a shortest $(w,x)$-path, whence $d(v,x)=k$. Notice also that since $w\in S(X^{\prime})\cap N(u)$, $w$ also belongs to $S(X)$. Thus, the entry $d(v,x)$ of $D^{\prime}$ is derived from the entry $d(w,x)$ of $D$.
Now suppose that $d(w,x)\leq k$ for every vertex $w\in S(X^{\prime})\cap N(u)$. We assert that in this case $d(v,x)=k-1$. Indeed, suppose it is not the case. Then necessarily, we must have $d(x,v)=k$. Let $K$ be a maximal simplex of the sphere $S_{k}(x)$ containing $u$ and $v$. Consider a maximal simplex $K^{\prime}$ of $X$ containing $K$. Both $K$ and $K^{\prime}$ are cliques of $G$. Since $u$ is a local maxima of the distance function $d(x,\cdot)$ and $K$ is a maximal clique of $S_{k}(x)$, we get $K^{\prime}\setminus K\subseteq S_{k-1}(x)$. On the other hand, by Lemma 2.2, $K^{\prime}\setminus K\neq\varnothing$. We claim that $K$ is contained in $\partial X\subseteq S(X)$. Indeed, let $K^{\prime\prime}=K^{\prime}\setminus\{y\}$ for some $y\in K^{\prime}\setminus K$ ($y$ exists since $K^{\prime}\setminus K\neq\varnothing$). If there was some simplex $K^{\prime\prime\prime}$ different from $K^{\prime}$ and strictly containing $K^{\prime\prime}$, then there would exist some $t$ adjacent in $G$ to all vertices of $K^{\prime\prime}$ and not belonging to $K^{\prime}\setminus K$. By the maximality choice of $K$ in $S_{k}(x)$, such $t$ would have to belong to the sphere $S_{k-1}(x)$. So $t$ would be adjacent to $y$, and hence to all vertices of $K^{\prime}$, a contradiction with the maximality choice of $K^{\prime}$. This contradiction shows that $v\in K\subseteq\partial X\subseteq S(X)$. Hence, $v$ would belong to $\partial X$, which is impossible. This proves that $d(v,x)=k-1$ and thus that the entry $d(v,x)$ of $D^{\prime}$ can be derived from the entry $d(u,x)$ of $D$ and the conditions of Case 2.
@@ 4. Boundary rigidity of Helly complexes
Next, from $S(X)$ and the distance matrix $D$ of $S(X)$ in $G$, we define a set $S(X^{\prime})$ such that $\partial X^{\prime}\subseteq S(X^{\prime})\subseteq V^{\prime}=V\setminus\{u\}$ and we show how to compute the distance matrix of $S(X^{\prime})$ in $G$ (and thus in $G^{\prime}$). We distinguish the same two cases. The analysis of Case 1 is completely the same because it uses only Lemma 2.5 which holds for all dismantlable graphs. Now we consider Case 2, i.e., the case where there does not exist a vertex $v^{\prime}\in S(X)$, $v^{\prime}\sim u$ such that $N[u]\cap S(X)\subseteq N[v^{\prime}]$. In this case, we also set $X^{\prime}=X\setminus\{u\}$, and $S(X^{\prime})=S(X)\setminus\{u\}\cup\{v\}$. Pick any $x\in S(X^{\prime})$ and we want to derive the distance $d(v,x)$. If $x\in N(u)$, then $d(v,x)=1$ since $v$ dominates $u$. Otherwise, let $k=d(x,u)\geq 2$. If there exists $w\in S(X^{\prime})\cap N(u)$ with $d(x,w)=k+1$, then as in case of systolic complexes we can argue that $d(x,v)=k$. Now suppose that $d(w,x)\leq k$ for every vertex $w\in S(X^{\prime})\cap N(u)$. We assert that in this case $d(v,x)=k-1$. Applying Lemma 2.3 to $x$ and $u$, there exists a vertex $v^{\prime}\sim u$ at distance $k-1$ from $x$ and adjacent to all vertices $w\in N[u]\cap B_{k}(x)$. Since any vertex $w\in N[u]$ has distance at most $k$ from $x$, we conclude that $v^{\prime}$ dominates $u$ in $G$. Since $v$ also dominates $u$ and $v\notin S(X)$, necessarily $v^{\prime}=v$, yielding $d(v,x)=k-1$. This allows to compute the distance matrix $D^{\prime}$ of $S(X^{\prime})$. Finally, the proof of Claim 1 and the final application of the induction are the same as for systolic complexes. This establishes Theorem 1.3.
@@ 4.2. The complexity of the reconstruction
At each step where $S(X^{\prime})=S(X)\setminus\{u\}\cup\{v\}$, we should compute $d(x,v)$ for any vertex $x\in S(X)\setminus\{u\}$. We know that either $d(x,v)=d(x,u)$ or $d(x,v)=d(x,u)-1$, and that we are in the second case if and only if $d(x,u)=\max\{d(x,w):w\in S(X)\text{ and }w\sim u\}$. Consequently, for every vertex $w$ in $S(X)$, we can compute $d(x,v)$ in $O(\deg(u))$ time and we can thus update the distance matrix $D$ as well as the graphs $G^{\prime}$ and $G[S(X)]$ in $O(n_{0}\cdot\deg(u))$ time. Therefore, for any Helly complex $X$, starting from the boundary $\partial X$ of size $n_{0}$ and its distance matrix, one can reconstruct $X$ in $O(m\cdot n_{0})$ time.
@@ 5.1. The computation of a dominated vertex
Suppose now that $|S_{1}|\geq 2$. Pick any $z_{2}\in S_{1}$ and consider a vertex $u\in S_{1}$ such that $u$ has a neighbor $v$ in $S_{1}$ with $d(z_{2},v)=d(z_{2},u)-1$ and $u$ has no neighbor $w$ in $S_{1}$ with $d(z_{2},w)=d(z_{2},u)+1$. Observe that we can find such a vertex $u$ among the vertices at maximum distance from $z_{2}$ in the connected component $S_{1}$. We want to show that $u$ is a dominated vertex. Let $k_{2}=d(z_{2},u)$. For that we need to consider two cases depending on whether $u$ has neighbors in $S_{1}$ at distance $k_{2}$ from $z$ or not.
@@
We claim that in this case, $v$ is the only neighbor of $u$ in $S_{1}$. Indeed, suppose that there exists $w\in N(u)\setminus\{w\}$ such that $d(z_{1},w)=k_{1}$. Since $d(z_{2},w)=k_{2}-1$, by Lemma 2.2, we get a triangle $uvw$ equidistant from $z_{1}$, a contradiction.
@@
For any $w\in N(u)\cap S_{1}$ such that $d(z_{2},w)=k_{2}$, by (TC), there is a vertex $x\sim u,w$ such that $d(z_{2},x)=k_{2}-1$. Note that $d(z_{1},x)=k_{1}-1$ as otherwise $uxw$ is a triangle equidistant from $z_{1}$, which is impossible. This implies in particular that $x\neq v$.

### v2

@@ Acknowledgements
The work presented here was partially supported by the Carlsberg Foundation, grant CF23-1225. J.C. and V.C. were also partially supported by the ANR project MIMETIQUE “Mineurs métriques” and J.C. was partially supported by the ANR project DUCAT (ANR-20-CE48-0006).
@@ 2.1. Graphs and simplicial complexes
All graphs $G=(V(G),E(G))$ considered in this paper are finite, undirected, connected, and contain neither multiple edges, nor loops. For two distinct vertices $v,w\in V(G)$ we write $v\sim w$ when there is an edge connecting $v$ with $w$. The subgraph of $G$ induced by a subset $A\subseteq V(G)$ is the graph $G[A]=(A,E^{\prime})$ such that $uv\in E^{\prime}$ if and only if $uv\in E(G)$. The distance $d(u,v)=d_{G}(u,v)$ between two vertices $u$ and $v$ of a graph $G$ is the length of a shortest (u,v)(u,v)–path. An induced subgraph $H=G[A]$ of a graph $G$ is an isometric subgraph of $G$ if $d_{H}(u,v)=d_{G}(u,v)$ for any two vertices $u,v\in A$. The interval $I(u,v)$ between $u$ and $v$ consists of all vertices on shortest (u,v)(u,v)–paths, that is, of all vertices (metrically) between $u$ and $v$: $I(u,v)=\{x\in V(G):d(u,x)+d(x,v)=d(u,v)\}$. An induced subgraph of $G$ is called convex if it includes the interval of $G$ between any pair of its vertices. For a vertex $v$ of $G$ and an integer $r\geq 1$, we will denote by $B_{r}(v)$ the ball in $G$ (and the subgraph induced by this ball) of radius $r$ centered at $v$, i.e., $B_{r}(v)=\{x\in V(G):d(v,x)\leq r\}.$ More generally, the rrball around a set $A\subseteq V(G)$ is the set (or the subgraph induced by) $B_{r}(A)=\{v\in V(G):d(v,A)\leq r\},$ where $d(v,A)=\mbox{min}\{d(v,x):x\in A\}$. Similarly, the sphere in $G$ centered at $v$ of radius $r$ is the set (or the subgraph induced by) $S_{r}(v)=\{x\in V(G):d(v,x)=r\}$. As usual, $N(v)=S_{1}(v)$ denotes the set of neighbors of a vertex $v$ in $G$, and $N[v]=B_{1}(v)$. The link of $v\in V(G)$ is the subgraph of $G$ induced by $N(v)$.
@@
An abstract simplicial complex $X$ on a finite set $V$ is a set of nonempty subsets of $V$, called simplices, such that any nonempty subset of a simplex is also a simplex. The dimension of a simplex $\sigma$ is $|\sigma|-1$ and the dimension of $X$ is the largest dimension of a simplex of $X$. If $\sigma$ is a simplex of dimension $k$, then all simplices of $X$ contained in $\sigma$ and having dimension k1k-1 are called the facets of $\sigma$. Simplices of $X$ are called maximal if they are maximal by inclusion, and non-maximal otherwise. The 2-dimensional simplices of $X$ are called triangles, and if a triangle is spanned by vertices $u$, $v$ and $w$, then we denote it by uvwuvw.
@@
For an abstract simplicial complex $X$, the kk–skeleton $X^{(k)}$ of $X$ consists of all simplices of dimension at most $k$. We use the notations $V(X)=X^{(0)}$ for the set of vertices (0-simplices) of $X$ and $G(X):=X^{(1)}$ for the 1-skeleton of $X$; the graph $G(X)$ will be always endowed with the standard graph distance $d_{G}$. The clique complex of a graph $G$ is the abstract simplicial complex having the cliques (i.e., complete subgraphs) of $G$ as simplices. A simplicial complex $X$ is a flag simplicial complex if $X$ is the clique complex of its 11–skeleton. Note that flag complexes are uniquely determined by their 1-skeleton.
@@
Each abstract simplicial complex $X$ gives rise to a geometric simplicial complex $X||$ obtained by replacing each simplex $\sigma$ of $X$ by an Euclidean simplex ${\sigma}||$ such that gluing maps between simplices are isometries (see the book of Hatcher [22]). Vice-versa, the set of simplices of a geometric simplicial complex define an abstract simplicial complex. Although the definitions of boundary and boundary rigidity presented in the next subsection are motivated by geometric simplicial complexes, all these notions can be defined in equivalent way for abstract simplicial complexes. Therefore, in the rest of the paper by a “simplicial complex” we will mean a finite abstract simplicial complex.
@@ 2.3. Bridged and Helly graphs
    Quadrangle Condition: for $v,x,y,u\in V$ with d(v,x)=d(v,y)=k=d(v,u)1d(v,x)=d(v,y)=k=d(v,u)-1 and $u\sim x,y$, $x\nsim y$, $\exists z\in B_{k-1}(v)$, $z\sim x,y$.
@@
Let $G$ be a Helly graph, and $z$ and $u$ be two vertices of $G$ with $k=d(z,u)\geq 2$. Then there exists a vertex $v$ which has distance k1k-1 to $z$ and is adjacent to $u$ and to all neighbors $w$ of $u$ with $d(z,w)\leq k$.
@@ 3. Boundary rigidity of systolic complexes
In this case, we set $X^{\prime}=X\setminus\{u\}$, and $S(X^{\prime})=S(X)\setminus\{u\}\cup\{v\}$. To derive the distance matrix $D^{\prime}$ of $S(X^{\prime})$ from the distance matrix $D$ of $S(X)$, we need to compute the distances $d(v,x)$ from $v$ to all vertices $x$ of $S(X^{\prime})$. Pick any $x\in S(X^{\prime})$. If $x\in N(u)$, then $d(v,x)=1$ since $v$ dominates $u$. Otherwise, let $k=d(x,u)$. If there exists $w\in S(X^{\prime})\cap N(u)$ with $d(x,w)=k+1$, then we can set $d(x,v)=k$. Indeed, $u$ belongs to a shortest $(w,x)$-path $P$ of $G$. If $y$ is the neighbor of $u$ in $P$, then $v\sim w,y$ because $v$ dominates $u$ in $G$. Consequently, $v$ is on a shortest $(w,x)$-path, whence $d(v,x)=k$. Notice also that since $\in S(X^{\prime})\cap N(u), ww$ also belongs to $S(X)$. Thus, the entry $d(v,x)$ of $D^{\prime}$ is derived from the entry $d(w,x)$ of $D$.
Now suppose that $d(w,x)\leq k$ for every vertex $w\in S(X^{\prime})\cap N(u)$. We assert that in this case d(v,x)=k1d(v,x)=k-1. Indeed, suppose it is not the case. Then necessarily, we must have $d(x,v)=k$. Let $K$ be a maximal simplex of the sphere $S_{k}(x)$ containing $u$ and $v$. Consider a maximal simplex $K^{\prime}$ of $X$ containing $K$. Both $K$ and $K^{\prime}$ are cliques of $G$. Since $u$ is a local maxima of the distance function $d(x,\cdot)$ and $K$ is a maximal clique of $S_{k}(x)$, we get $K^{\prime}\setminus K\subseteq S_{k-1}(x)$. On the other hand, by Lemma 2.2, $K^{\prime}\setminus K\neq\varnothing$. We claim that $K$ is contained in $\partial X\subseteq S(X)$. Indeed, let $K^{\prime\prime}=K^{\prime}\setminus\{y\}$ for some $y\in K^{\prime}\setminus K$ ($y$ exists since $K^{\prime}\setminus K\neq\varnothing$). If there was some simplex $K^{\prime\prime\prime}$ different from $K^{\prime}$ and strictly containing $K^{\prime\prime}$, then there would exist some $t$ adjacent in $G$ to all vertices of $K^{\prime\prime}$ and not belonging to $K^{\prime}\setminus K$. By the maximality choice of $K$ in $S_{k}(x)$, such $t$ would have to belong to the sphere $S_{k-1}(x)$. So $t$ would be adjacent to $y$, and hence to all vertices of $K^{\prime}$, a contradiction with the maximality choice of $K^{\prime}$. This contradiction shows that $v\in K\subseteq\partial X\subseteq S(X)$. Hence, $v$ would belong to $\partial X$, which is impossible. This proves that d(v,x)=k1d(v,x)=k-1 and thus that the entry $d(v,x)$ of $D^{\prime}$ can be derived from the entry $d(u,x)$ of $D$ and the conditions of Case 2.
@@ 4. Boundary rigidity of Helly complexes
Next, from $S(X)$ and the distance matrix $D$ of $S(X)$ in $G$, we define a set $S(X^{\prime})$ such that $\partial X^{\prime}\subseteq S(X^{\prime})\subseteq V^{\prime}=V\setminus\{u\}$ and we show how to compute the distance matrix of $S(X^{\prime})$ in $G$ (and thus in $G^{\prime}$). We distinguish the same two cases. The analysis of Case 1 is completely the same because it uses only Lemma 2.5 which holds for all dismantlable graphs. Now we consider Case 2, i.e., the case where there does not exist a vertex $v^{\prime}\in S(X)$, $v^{\prime}\sim u$ such that $N[u]\cap S(X)\subseteq N[v^{\prime}]$. In this case, we also set $X^{\prime}=X\setminus\{u\}$, and $S(X^{\prime})=S(X)\setminus\{u\}\cup\{v\}$. Pick any $x\in S(X^{\prime})$ and we want to derive the distance $d(v,x)$. If $x\in N(u)$, then $d(v,x)=1$ since $v$ dominates $u$. Otherwise, let $k=d(x,u)\geq 2$. If there exists $w\in S(X^{\prime})\cap N(u)$ with $d(x,w)=k+1$, then as in case of systolic complexes we can argue that $d(x,v)=k$. Now suppose that $d(w,x)\leq k$ for every vertex $w\in S(X^{\prime})\cap N(u)$. We assert that in this case d(v,x)=k1d(v,x)=k-1. Applying Lemma 2.3 to $x$ and $u$, there exists a vertex $v^{\prime}\sim u$ at distance k1k-1 from $x$ and adjacent to all vertices $w\in N[u]\cap B_{k}(x)$. Since any vertex $w\in N[u]$ has distance at most $k$ from $x$, we conclude that $v^{\prime}$ dominates $u$ in $G$. Since $v$ also dominates $u$ and $v\notin S(X)$, necessarily $v^{\prime}=v$, yielding d(v,x)=k1d(v,x)=k-1. This allows to compute the distance matrix $D^{\prime}$ of $S(X^{\prime})$. Finally, the proof of Claim 1 and the final application of the induction are the same as for systolic complexes. This establishes Theorem 1.3.
@@ 4.2. The complexity of the reconstruction
At each step where $S(X^{\prime})=S(X)\setminus\{u\}\cup\{v\}$, we should compute $d(x,v)$ for any vertex $x\in S(X)\setminus\{u\}$. We know that either $d(x,v)=d(x,u)$ or d(x,v)=d(x,u)1d(x,v)=d(x,u)-1, and that we are in the second case if and only if $d(x,u)=\max\{d(x,w):w\in S(X)\text{ and }w\sim u\}$. Consequently, for every vertex $w$ in $S(X)$, we can compute $d(x,v)$ in $O(\deg(u))$ time and we can thus update the distance matrix $D$ as well as the graphs $G^{\prime}$ and $G[S(X)]$ in $O(n_{0}\cdot\deg(u))$ time. Therefore, for any Helly complex $X$, starting from the boundary $\partial X$ of size $n_{0}$ and its distance matrix, one can reconstruct $X$ in $O(m\cdot n_{0})$ time.
@@ 5.1. The computation of a dominated vertex
Suppose now that $|S_{1}|\geq 2$. Pick any $z_{2}\in S_{1}$ and consider a vertex $u\in S_{1}$ such that $u$ has a neighbor $v$ in $S_{1}$ with $d(z_{2},v)=d(z_{2},u)-1$ and $u$ has no neighbor $w$ in $S_{1}$ with $d(z_{2},w)=d(z_{2},u)+1$. Observe that we can find such a vertex $u$ among the vertices at maximum distance from $z_{2}$ in the connected component $S_{1}$. We want to show that $u$ is a dominated vertex. Let $k_{2}=d(z_{2},u)$. For that we need to consider two cases depending on whether $u$ has neighbors in $S_{1}$ at distance $k_{2}$ from $z_{2}$ or not.
@@
We claim that in this case, $v$ is the only neighbor of $u$ in $S_{1}$. Indeed, suppose that there exists $w\in N(u)\setminus\{v\}$ such that $d(z_{1},w)=k_{1}$. Since $d(z_{2},w)=k_{2}-1$, by Lemma 2.2, we get a triangle uvwuvw equidistant from $z_{1}$, a contradiction.
@@
For any $w\in N(u)\cap S_{1}$ such that $d(z_{2},w)=k_{2}$, by (TC), there is a vertex $x\sim u,w$ such that $d(z_{2},x)=k_{2}-1$. Note that $d(z_{1},x)=k_{1}-1$ as otherwise uxwuxw is a triangle equidistant from $z_{1}$, which is impossible. This implies in particular that $x\neq v$.